# MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification

### v3

@@ 3.1 Bridging Pathology Visual and Text Encoders
Given a set of collected pathology image-text pairs $\left\{(\mathbf{I}_{i},\mathbf{T}_{i})|\,i=1,2..,N\,\right\}$ (Sec. 4), we denote by $E_{I}(.)$ be a pre-trained vision encoder from Prov-GigaPath, extracting patch-level feature, and $E_{T}(.)$ the pre-trained text encoder from PLIP model. Given a batch size of $B$ samples, the image and text embeddings are computed as $\mathbf{x}_{i}=E_{I}(\mathbf{I}_{i})\in\mathbb{R}^{d_{v}},\,\mathbf{t}_{i}=E_{T}(\mathbf{T}_{i})\in\mathbb{R}^{d_{t}}$. We then design two trainable adaptors $A_{I}(.)$ and $A_{T}(.)$, that maps $\left(\mathbf{x}_{i},\mathbf{t}_{i}\right)$ into the same hidden dimension $\mathbb{R}^{d}$ and minimizes the noise contrastive loss Oord et al. (2018):
@@
where $\cos(.)$ is the cosine similarity, and $\tau$ denotes for temperature of the softmax function. For parameter efficiency, we train only the adaptors $A_{I}(.),A_{T}(.)$ while keeping the Prov-GigaPath visual encoder and PLIP text encoder frozen. After optimizing Eq.(1), we use the outputs of the adaptors as visual and text embeddings for downstream tasks. Unless otherwise specified, we refer to this model as GigaPath-PLIP.
@@ 3.2 Multi-Magnification Descriptive Text Prompts
To improve vision-language models (VLMs) for whole-slide image (WSI) analysis, designing effective text prompts is essential. Pathologists typically examine WSIs by first assessing tissue structures at low magnification before zooming in to analyze finer details such as nuclear size and shape. Inspired by this diagnostic workflow and the inherently multi-scale nature of WSIs, recent studies  Shi et al. (2024); Han et al. (2024) have introduced dual-scale visual descriptive text prompts to guide VLMs, leading to significant improvements in classification performance. Building on this observation, we further extend and refine this strategy to enhance model effectiveness.
@@
Second, at each low/high scale, rather than inserting a single learnable text prompt of length $K$ alongside a frozen contextual prompt from LLMs Shi et al. (2024); Han et al. (2024), we propose using $M$ learnable prompts. This approach aims to capture different sub-regions or structural features within each patch that might be overlooked with only a single prompt. Specifically, we define visual descriptive text prompts for both low and high resolutions as follows:
@@
where $[\omega_{i}^{\beta}]_{j},j\in[1,...,K],i\in[1,..,M]$ are $KM$ trainable textual prompts for each resolution $\beta\in\{l,h\}$.
@@ 3.3 Granularity-aware Visual Prompt Learning
We propose to adapt visual prompts to frozen features extracted by a pre-trained vision encoder in the VLM model by taking into account the image patch level and spatial groupings of patches. Specifically, for each WSI $W$, we denote by $\left\{W^{(l)},W^{(h)}\right\}$ are representations of $W$ at low and high magnification. We define a bag of multiple instances of $W$ as $I=\left\{I^{(l)},I^{(h)}\right\}$ where $I^{(l)}\in\mathbb{R}^{N_{l}\times N_{b}\times N_{b}\times 3}$, $I^{(h)}\in\mathbb{R}^{N_{h}\times N_{b}\times N_{b}\times 3}$ with $N_{l},N_{h}$ indicate the number of low and high-resolution image patches and $N_{b}$ is the patch size. Following prior works Shi et al. (2024); Ilse et al. (2018); Lu et al. (2021); Han et al. (2024), we employ a non-overlapping sliding window technique to extract patches $I$ from the WSI.
@@ 3.3.1 Patches-based Prompting
The frozen image encoder $E_{I}(.)$ (or $A_{I}(E_{I}(.))$ in case of GigaPath-PLIP) is used to map patches $I$ into a feature vector $H=\{H^{(l)}\in\mathbb{R}^{N_{l}\times d},H^{(h)}\in\mathbb{R}^{N_{h}\times d}\}$ where $d$ denotes the feature dimension. To effectively consolidate the extensive set of patch features into a final slide-level representation, we introduce a set of learnable visual prompts $\mathbf{p}_{v}\in\mathbb{R}^{N_{p}\times d}$, which facilitate the progressive merging of patch features in $H^{(l)}$(similarly for $H^{(h)}$) (Figure 2). In particular, we formulate $\mathbf{p}_{v}$ as Query and take all features in $H^{(l)}$ as the Keys $K^{(l)}_{p}$ and Values $V^{(l)}_{p}$ in in self-attention Vaswani (2017). We then associate $\mathbf{p}_{v}$ with patch features as:
@@
where $\operatorname{Normalize(.)}$ and $\operatorname{Softmax(.)}$ indicate the layer normalization operator and activation function respectively. Intuitively, Eq.(3) computes the correlations between the visual prompt $\mathbf{p}_{v}$ and all individual feature patches in $H^{l}$, subsequently grouping patches with high similarity to form fused prompt embeddings. However, since cancerous tissues in WSIs often appear as large, contiguous regions of adjacent image patches, this motivates the introduction of spatial patch group-based attention.
@@ 3.3.2 Spatial Patch Group-based Prompting
We next design a trainable message-passing network $g_{\epsilon}(.)$ based on the graph attention layer (GAT) Veličković et al. (2017) to capture the feature representation of each node and its local neighbors. The message passing of the GAT layer is formulated as:
@@
After doing a message passing by $g_{\epsilon}(.)$, the graph of patch-image features $G^{(l)}$ is updated to $G^{(l)^{\prime}}$, where each node now represents a super-node that encapsulates its corresponding feature region. We then squeeze all feature nodes in $G^{(l)^{\prime}}$ as a vector $H^{(l)}_{gr}$ and treat them as another Keys $K_{gr}^{(l)}$ and Values $V_{gr}^{(l)}$ for region-level features. Similar to Eq.(3), we associate prompt $\mathbf{p}_{v}$ with those group-level features:
@@ 3.4 Optimal Transport for Visual-Text Alignment
Given descriptive text prompts $\mathbf{T}^{(l)}$ and $\mathbf{T}^{(h)}$ (Eq.(2)) and visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}$ and $\mathbf{p}_{v}^{(h)}$ (Eq.(6)) for low and high resolutions, our goal is to maximize the similarity between slide and text embeddings for each class $c$. Rather than relying on cosine distance, as in prior works Zhou et al. (2022b, a); Zhao et al. (2024); Qu et al. (2024a, a); Singh and Jaggi (2020), we propose using optimal transport (OT)-based distance to capture a more nuanced cross-alignment between visual and text domains. Although OT has been explored for prompt learning in natural images and multi-modal learning Kim et al. (2023); Chen et al. (2023); Nguyen et al. (2024a); Séjourné et al. (2023), we are the first to adapt it for whole-slide imaging (WSI), effectively handling the alignment of multi-magnification patches to capture rich structural details across scales.
@@
$\bm{\mu}=\sum^{M}_{i=1}p_{i}\delta_{f_{i}},\quad\bm{\nu}=\sum^{N}_{j=1}q_{j}\delta_{g_{i}},$ (7)
where $\delta_{f}$ and $\delta_{g}$ represent Dirac delta functions centered at $\bm{f}$ and $\bm{g}$, respectively, and $M$ and $N$ indicate the dimensions of the empirical distribution. The weight vectors $\bm{p}=\{p_{i}\}^{M}_{i=1}$ and $\bm{q}=\{q_{i}\}^{N}_{j=1}$ lie within the $M$ and $N$-dimensional simplex, respectively, meaning they satisfy $\sum_{i=1}^{M}p_{i}=1$ and $\sum_{j=1}^{N}q_{j}=1$. The discrete optimal transport problem can then be expressed as:
$\bm{T}^{\ast}=\underset{\bm{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\bm{T}_{ij}\bm{C}_{ij}$
[…]
where $\bm{T}^{\ast}$ is denoted as the optimal transport plan, which is optimized to minimize the total distance between the two probability vectors, $\bm{C}$ is the cost matrix which measures the distance between $\bm{f}_{i}$ and $\bm{g}_{j}$. We then define the OT distance between $\bm{\mu}$ and $\bm{\nu}$ as:
$d_{\mathrm{OT}}(\bm{\mu},\bm{\nu})=\langle\bm{T}^{\ast},\bm{C}\rangle.$ (9)
@@
Given the visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}\in\mathbb{R}^{N_{p}\times d}$ in Eq.(6) and the descriptive text prompts $\mathbf{T}^{(l)}$ in Eq.(2), we compute the textual embedding for $\mathbf{T}^{(l)}$ as $\mathbf{p}_{t}^{(l)}=E_{T}(\mathbf{T}^{(l)})\in\mathbb{R}^{M\times d}$. We next denote $\mathbf{T}^{(l)}_{c}$ as the input text prompts, $\left(\mathbf{p}_{t}^{(l)}\right)_{c}$ as the extracted textual embedding, and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$ as the visual prompt-guided slide features associated with class $c$. We then aim to minimize the distance between $\mathbf{T}^{(l)}_{c}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$, indicated as $d_{\mathrm{OT}}\left(\mathbf{T}^{(l)}_{c},\left(\mathbf{p}_{v}^{(l)}\right)_{c}\right)$ in the paper, by computing optimal transport distance between $\left(\mathbf{p}_{t}^{(l)}\right)_{c}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$. Specifically, we treat $\left(\mathbf{p}_{t}^{(l)}\right)_{c}\rightarrow\bm{F}=\left\{\bm{f}_{i}|_{i=1}^{M}\right\}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}\rightarrow\bm{G}=\left\{\bm{g}_{j}|_{j=1}^{N_{p}}\right\}$ and compute the cost matrix $\bm{C}$ as   $\bm{C}=\left(\bm{1}-\bm{F}^{T}\,\bm{G}\right)\in\mathbb{R}^{M\times N_{p}}$, which used to compute $\bm{T}^{\ast}$ in Eq. (8) for estimate optimal transport distance defined in Eq. (9). Following the same procedure, we can also compute $d_{\mathrm{OT}}\left(\mathbf{T}^{(h)}_{c},\left(\mathbf{p}_{v}^{(h)}\right)_{c}\right)$ at high-resolution image patches. Then, the prediction probability is written as:
@@
The details for solvers of Eq.(9) and a relaxed version with unbalanced optimal transport are presented in Sections (D) and (D.1) in Appendix. Intuitively, using OT, in this case, offers several key advantages over cosine similarity. Pathology images exhibit complex, heterogeneous patterns that can be described from multiple perspectives. OT models these relationships as a distribution, enabling a more holistic alignment that handles variability and incomplete details while reducing noise from irrelevant prompts. This enhances the model’s ability to generalize to unseen or complex disease cases.
@@ 4.1 Settings
PatchGastricADC22Tsuneki and Kanavati (2022) consists of approximately 262K patches derived from WSI of H&E-stained gastric adenocarcinoma specimens, each paired with associated diagnostic captions collected from the University of Health and Welfare, Mita Hospital, Japan. QUILT-1M Ikezogwo et al. (2024) includes approximately 653K images and one million pathology image-text pairs, gathered from 1,087 hours of educational histopathology videos presented by pathologists on YouTube. ARCH Gamper and Rajpoot (2021) is a pathology multiple-instance captioning dataset containing pathology images at the bag and tile level. However, our work focuses on tile-level images from all datasets for our contrastive training strategy. In total, we collected approximately 923K images from these datasets.
@@
For the classification task, the proposed method was evaluated in three datasets from the Cancer Genome Atlas Data PortalThe Cancer Genome Atlas () (TCGA): TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA. We followed the ViLa-MILShi et al. (2024) experimental settings for TCGA-NSCLC and TCGA-RCC, randomly selecting proportions for training, validation, and testing. For TCGA-BRCA, we adapted the training and testing slide ID from MSCPT Han et al. (2024). The detailed description is included in the appendix section.
@@
Table 1: Comparison of methods on TCGA-BRCA with few-shot settings. Results are shown for AUC, F1, and Accuracy (ACC). FVM denotes for foundation vision-language models.
CLIP ImageNet Pretrained Methods # Param. TCGA-BRCA
AUC F1 ACC
Max-pooling 197K 60.42$\pm$4.35 56.40$\pm$3.58 68.55$\pm$6.54
Mean-pooling 197K 66.64$\pm$4.21 60.70$\pm$2.78 71.73$\pm$3.59
ABMIL Ilse et al. (2018) 461K 69.24$\pm$3.90 61.72$\pm$3.36 72.77$\pm$3.15
CLAM-SB Lu et al. (2021) 660K 67.80$\pm$5.14 60.51$\pm$5.01 72.46$\pm$4.36
CLAM-MB Lu et al. (2021) 660K 60.81$\pm$4.87 55.48$\pm$4.96 67.31$\pm$4.19
TransMIL Shao et al. (2021) 2.54M 65.62$\pm$3.20 60.75$\pm$4.04 67.52$\pm$4.16
DSMIL Li et al. (2021) 462K 66.18$\pm$3.08 59.35$\pm$3.18 67.52$\pm$1.56
RRT-MIL Tang et al. (2024) 2.63M 66.33$\pm$4.30 61.14$\pm$5.93 71.21$\pm$8.94
CoOp Zhou et al. (2022b) 337K 68.86$\pm$4.35 61.64$\pm$2.40 71.08$\pm$3.22
CoCoOp Zhou et al. (2022a) 370K 69.13$\pm$4.27 61.48$\pm$2.62 72.41$\pm$1.87
Metaprompt Zhao et al. (2024) 360K 69.12$\pm$4.46 63.39$\pm$4.38 74.65$\pm$7.20
TOP Qu et al. (2024a) 2.11M 69.74$\pm$3.14 63.39$\pm$4.62 74.41$\pm$5.27
ViLa-MIL Shi et al. (2024) 2.77M 72.25$\pm$6.16 62.04$\pm$2.38 75.01$\pm$6.14
MSCPT Han et al. (2024) 1.35M 74.56$\pm$4.54 65.59$\pm$1.85 75.82$\pm$2.38
MGPath (ViT) 592K 74.96$\pm$6.98 64.60$\pm$5.39 77.10$\pm$2.39
FVM CONCH Lu et al. (2024) 110M 84.11$\pm$15.44 65.63$\pm$10.81 73.24$\pm$8.89
QUILT  Ikezogwo et al. (2024) 63M 73.48$\pm$10.57 63.78$\pm$8.72 73.26$\pm$10.13
PLIP Pathology Pretrained Max-pooling 197K 66.50$\pm$2.74 61.50$\pm$2.88 71.57$\pm$4.82
Mean-pooling 197K 71.62$\pm$2.41 66.34$\pm$2.96 74.45$\pm$2.49
ABMIL Ilse et al. (2018) 461K 72.41$\pm$4.25 63.04$\pm$3.62 74.09$\pm$4.38
CLAM-SB Lu et al. (2021) 660K 72.34$\pm$6.17 65.51$\pm$3.28 76.16$\pm$4.36
CLAM-MB Lu et al. (2021) 660K 73.41$\pm$3.76 66.11$\pm$1.94 77.88$\pm$2.30
TransMIL Shao et al. (2021) 2.54M 74.98$\pm$6.01 67.50$\pm$6.00 77.04$\pm$6.14
DSMIL Li et al. (2021) 462K 71.44$\pm$2.72 64.48$\pm$1.64 75.26$\pm$2.28
RRT-MIL Tang et al. (2024) 2.63M 71.21$\pm$6.46 64.15$\pm$1.38 75.92$\pm$5.10
CoOp Zhou et al. (2022b) 337K 71.53$\pm$2.45 64.84$\pm$2.40 74.22$\pm$5.02
CoCoOp Zhou et al. (2022a) 370K 72.65$\pm$4.63 66.63$\pm$3.55 66.98$\pm$3.35
Metaprompt Zhao et al. (2024) 360K 74.86$\pm$4.25 65.03$\pm$1.81 77.88$\pm$3.22
TOP Qu et al. (2024a) 2.11M 76.13$\pm$6.01 66.55$\pm$1.72 78.58$\pm$5.30
ViLa-MIL Shi et al. (2024) 2.77M 74.06$\pm$4.62 66.03$\pm$1.81 78.12$\pm$4.88
MSCPT Han et al. (2024) 1.35M 75.55$\pm$5.25 67.46$\pm$2.43 79.14$\pm$2.63
MGPath 592K 79.02$\pm$6.43 68.25$\pm$4.42 79.65$\pm$1.72
MGPath (PLIP-G) 5.35M 87.36$\pm$1.85 73.13$\pm$3.49 79.56$\pm$4.77
To implement spatial attention, we use a Graph Attention Network (GAT) to model spatial relationships between WSI patches. Each tile-level embedding serves as a node, connected to its left, right, top, and bottom neighbors, ensuring local spatial dependencies are captured. We then integrate spatial patch group-based attention $\mathbf{p}_{v,gr}$ into patch-based attention $\mathbf{p}_{v,p}$ using Equation 6. The hyperparameter $\alpha$ (0 to 1) controls the balance between spatial context and prototype-based guidance.

### v5

@@ 3.1 Bridging Pathology Visual and Text Encoders
Given a set of collected pathology image-text pairs $\left\{(\mathbf{I}_{i},\mathbf{T}_{i})|\,i=1,2..,N\,\right\}$ (Sec. 4), we denote by $E_{I}(.)$ be a pre-trained vision encoder from Prov-GigaPath, extracting patch-level feature, and $E_{T}(.)$ the pre-trained text encoder from PLIP model. Given a batch size of $B$ samples, the image and text embeddings are computed as $\mathbf{x}_{i}=E_{I}(\mathbf{I}_{i})\in\mathbb{R}^{d_{v}},\,\mathbf{t}_{i}=E_{T}(\mathbf{T}_{i})\in\mathbb{R}^{d_{t}}$. We then design two trainable adaptors $A_{I}(.)$ and $A_{T}(.)$, that maps $\left(\mathbf{x}_{i},\mathbf{t}_{i}\right)$ into the same hidden dimension $\mathbb{R}^{d}$ and minimizes the noise contrastive loss (Oord et al., 2018):
@@
where $\cos(.)$ is the cosine similarity, and $\tau$ denotes for temperature of the softmax function. For parameter efficiency, we train only the adaptors $A_{I}(.),A_{T}(.)$ while keeping the Prov-GigaPath visual encoder and PLIP text encoder frozen. After optimizing (1), we use the outputs of the adaptors as visual and text embeddings for downstream tasks. Unless otherwise specified, we refer to this model as MGPATH(PLIP-G).
@@ 3.2 Multi-Magnification Descriptive Text Prompts
To improve vision-language models (VLMs) for whole-slide image (WSI) analysis, designing effective text prompts is essential. Pathologists typically examine WSIs by first assessing tissue structures at low magnification before zooming in to analyze finer details such as nuclear size and shape. Inspired by this diagnostic workflow and the inherently multi-scale nature of WSIs, recent studies  (Shi et al., 2024; Han et al., 2024) have introduced dual-scale visual descriptive text prompts to guide VLMs, leading to significant improvements in classification performance. Building on this observation, we further extend and refine this strategy to enhance model effectiveness.
@@
Second, at each low/high scale, rather than inserting a single learnable text prompt of length $K$ alongside a frozen contextual prompt from LLMs (Shi et al., 2024; Han et al., 2024), we propose using $M$ learnable prompts. This approach aims to capture different sub-regions or structural features within each patch that might be overlooked with only a single prompt. Specifically, we define visual descriptive text prompts for both low and high resolutions as follows:
@@
where $\omega_{i}^{\beta}]_{j},j\in[1,...,K],i\in[1,..,M]$ are $KM$ trainable textual prompts for each resolution $\beta\in\{l,h\}$.
@@ 3.3 Granularity-aware Visual Prompt Learning
We propose to adapt visual prompts to frozen features extracted by a pre-trained vision encoder in the VLM model by taking into account the image patch level and spatial groupings of patches. Specifically, for each WSI $W$, we denote by $\left\{W^{(l)},W^{(h)}\right\}$ are representations of $W$ at low and high magnification. We define a bag of multiple instances of $W$ as $I=\left\{I^{(l)},I^{(h)}\right\}$ where $I^{(l)}\in\mathbb{R}^{N_{l}\times N_{b}\times N_{b}\times 3}$, $I^{(h)}\in\mathbb{R}^{N_{h}\times N_{b}\times N_{b}\times 3}$ with $N_{l},N_{h}$ indicate the number of low and high-resolution image patches and $N_{b}$ is the patch size. Following prior works (Shi et al., 2024; Ilse et al., 2018; Lu et al., 2021; Han et al., 2024), we employ a non-overlapping sliding window technique to extract patches $I$ from the WSI.
@@ 3.3.1 Patches-based Prompting
The frozen image encoder $E_{I}(.)$ (or $A_{I}(E_{I}(.))$ in case of MGPATH(PLIP-G)) is used to map patches $I$ into a feature vector $H=\{H^{(l)}\in\mathbb{R}^{N_{l}\times d},H^{(h)}\in\mathbb{R}^{N_{h}\times d}\}$ where $d$ denotes the feature dimension. To effectively consolidate the extensive set of patch features into a final slide-level representation, we introduce a set of learnable visual prompts $\mathbf{p}_{v}\in\mathbb{R}^{N_{p}\times d}$, which facilitate the progressive merging of patch features in $H^{(l)}$(similarly for $H^{(h)}$) (Figure 2). In particular, we formulate $\mathbf{p}_{v}$ as Query and take all features in $H^{(l)}$ as the Keys $K^{(l)}_{p}$ and Values $V^{(l)}_{p}$ in in self-attention (Vaswani, 2017). We then associate $\mathbf{p}_{v}$ with patch features as:
@@
where $\operatorname{Normalize(.)}$ and $\operatorname{Softmax(.)}$ indicate the layer normalization operator and activation function respectively. Intuitively, (3) computes the correlations between the visual prompt $\mathbf{p}_{v}$ and all individual feature patches in $H^{l}$, subsequently grouping patches with high similarity to form fused prompt embeddings. However, since cancerous tissues in WSIs often appear as large, contiguous regions of adjacent image patches, this motivates the introduction of spatial patch group-based attention.
@@ 3.3.2 Spatial Patch Group-based Prompting
We next design a trainable message-passing network $g_{\epsilon}(.)$ based on the graph attention layer (GAT) (Veličković et al., 2017) to capture the feature representation of each node and its local neighbors. The message passing of the GAT layer is formulated as:
@@
After doing a message passing by $g_{\epsilon}(.)$, the graph of patch-image features $G^{(l)}$ is updated to $G^{(l)^{\prime}}$, where each node now represents a super-node that encapsulates its corresponding feature region. We then squeeze all feature nodes in $G^{(l)^{\prime}}$ as a vector $H^{(l)}_{gr}$ and treat them as another Keys $K_{gr}^{(l)}$ and Values $V_{gr}^{(l)}$ for region-level features. Similar to (3), we associate prompt $\mathbf{p}_{v}$ with those group-level features:
@@ 3.4 Optimal Transport for Visual-Text Alignment
Given descriptive text prompts $\mathbf{T}^{(l)}$ and $\mathbf{T}^{(h)}$ ((2)) and visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}$ and $\mathbf{p}_{v}^{(h)}$ ((6)) for low and high resolutions, our goal is to maximize the similarity between slide and text embeddings for each class $c$. Rather than relying on cosine distance, as in prior works (Zhou et al., 2022b; a; Zhao et al., 2024; Qu et al., 2024a; a; Singh & Jaggi, 2020), we propose using optimal transport (OT)-based distance to capture a more nuanced cross-alignment between visual and text domains. Although OT has been explored for prompt learning in natural images and multi-modal learning (Kim et al., 2023; Chen et al., 2023; Nguyen et al., 2024a; Séjourné et al., 2023), we are the first to adapt it for whole-slide imaging (WSI), effectively handling the alignment of multi-magnification patches to capture rich structural details across scales.
@@
$\boldsymbol{\mu}=\sum^{M}_{i=1}p_{i}\delta_{f_{i}},\quad\boldsymbol{\nu}=\sum^{N}_{j=1}q_{j}\delta_{g_{i}},$ (7)
where $\delta_{f}$ and $\delta_{g}$ represent Dirac delta functions centered at $\boldsymbol{f}$ and $\boldsymbol{g}$, respectively, and $M$ and $N$ indicate the dimensions of the empirical distribution. The weight vectors $\boldsymbol{p}=\{p_{i}\}^{M}_{i=1}$ and $\boldsymbol{q}=\{q_{i}\}^{N}_{j=1}$ lie within the $M$ and $N$-dimensional simplex, respectively, meaning they satisfy $\sum_{i=1}^{M}p_{i}=1$ and $\sum_{j=1}^{N}q_{j}=1$. The discrete optimal transport problem can then be expressed as:
$\boldsymbol{T}^{\ast}=\underset{\boldsymbol{T}\in\mathbb{R}^{MXN}}{\arg{\min}}\sum^{M}_{i=1}\sum^{N}_{j=1}\boldsymbol{T}_{ij}\boldsymbol{C}_{ij}$
$\textrm{s.t.}\quad\boldsymbol{T}\boldsymbol{1}^{N}=\boldsymbol{\mu},\quad\boldsymbol{T}^{\top}\boldsymbol{1}^{M}=\boldsymbol{\nu}.$ (8)
where $\boldsymbol{T}^{\ast}$ is denoted as the optimal transport plan, which is optimized to minimize the total distance between the two probability vectors, $\boldsymbol{C}$ is the cost matrix which measures the distance between $\boldsymbol{f}_{i}$ and $\boldsymbol{g}_{j}$. We then define the OT distance between $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ as:
$d_{\mathrm{OT}}(\boldsymbol{\mu},\boldsymbol{\nu})=\langle\boldsymbol{T}^{\ast},\boldsymbol{C}\rangle.$ (9)
@@
Given the visual prompt-guided slide features $\mathbf{p}_{v}^{(l)}\in\mathbb{R}^{N_{p}\times d}$ in (6) and the descriptive text prompts $\mathbf{T}^{(l)}$ in (2), we compute the textual embedding for $\mathbf{T}^{(l)}$ as $\mathbf{p}_{t}^{(l)}=E_{T}(\mathbf{T}^{(l)})\in\mathbb{R}^{M\times d}$.
We next denote $\mathbf{T}^{(l)}_{c}$ as the input text prompts, $\left(\mathbf{p}_{t}^{(l)}\right)_{c}$ as the extracted textual embedding, and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$ as the visual prompt-guided slide features associated with class $c$. We then aim to minimize the distance between $\mathbf{T}^{(l)}_{c}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$, indicated as $d_{\mathrm{OT}}\left(\mathbf{T}^{(l)}_{c},\left(\mathbf{p}_{v}^{(l)}\right)_{c}\right)$ in the paper, by computing optimal transport distance between $\left(\mathbf{p}_{t}^{(l)}\right)_{c}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}$. Specifically, we treat $\left(\mathbf{p}_{t}^{(l)}\right)_{c}\rightarrow\boldsymbol{F}=\left\{\boldsymbol{f}_{i}|_{i=1}^{M}\right\}$ and $\left(\mathbf{p}_{v}^{(l)}\right)_{c}\rightarrow\boldsymbol{G}=\left\{\boldsymbol{g}_{j}|_{j=1}^{N_{p}}\right\}$ and compute the cost matrix $\boldsymbol{C}$ as   $\boldsymbol{C}=\left(\boldsymbol{1}-\boldsymbol{F}^{T}\,\boldsymbol{G}\right)\in\mathbb{R}^{M\times N_{p}}$, which used to compute $\boldsymbol{T}^{\ast}$ in (8) for estimate optimal transport distance defined in (9). Following the same procedure, we can also compute $d_{\mathrm{OT}}\left(\mathbf{T}^{(h)}_{c},\left(\mathbf{p}_{v}^{(h)}\right)_{c}\right)$ at high-resolution image patches. Then, the prediction probability is written as:
@@
The details for solvers of (9) and a relaxed version with unbalanced optimal transport are presented in Sections (G) and (G.1) in Appendix. Intuitively, using OT, in this case, offers several key advantages over cosine similarity. Pathology images exhibit complex, heterogeneous patterns that can be described from multiple perspectives. OT models these relationships as a distribution, enabling a more holistic alignment that handles variability and incomplete details while reducing noise from irrelevant prompts. This enhances the model’s ability to generalize to unseen or complex disease cases.
@@ 4.1 Settings
PatchGastricADC22 (Tsuneki & Kanavati, 2022) consists of approximately 262K patches derived from WSI of H&E-stained gastric adenocarcinoma specimens, each paired with associated diagnostic captions collected from the University of Health and Welfare, Mita Hospital, Japan. QUILT-1M (Ikezogwo et al., 2024) includes approximately 653K images and one million pathology image-text pairs, gathered from 1,087 hours of educational histopathology videos presented by pathologists on YouTube. ARCH (Gamper & Rajpoot, 2021) is a pathology multiple-instance captioning dataset containing pathology images at the bag and tile level. However, our work focuses on tile-level images from all datasets for our contrastive training strategy. In total, we collected approximately 923K images from these datasets.
@@
For the classification task, the proposed method was evaluated in three datasets from the Cancer Genome Atlas Data Portal (The Cancer Genome Atlas (), TCGA): TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA. We followed the ViLa-MIL(Shi et al., 2024) experimental settings for TCGA-NSCLC and TCGA-RCC, randomly selecting proportions for training, validation, and testing. For TCGA-BRCA, we adapted the training and testing slide ID from MSCPT (Han et al., 2024). The detailed description is included in the appendix section.
Evaluation Metrics.
For all experiments, we report the Area Under the Curve (AUC), F1 score (F1), and accuracy (ACC) as evaluation metrics. Each experiment is repeated five times, and we present the mean and standard deviation across these runs to ensure robustness and reproducibility.
@@
To implement spatial attention, we use a Graph Attention Network (GAT) to model spatial relationships between WSI patches. Each tile-level embedding serves as a node, connected to its left, right, top, and bottom neighbors, ensuring local spatial dependencies are captured. We then integrate spatial patch group-based attention $\mathbf{p}_{v,gr}$ into patch-based attention $\mathbf{p}_{v,p}$ using (6). The hyperparameter $\alpha$ (0 to 1) controls the balance between spatial context and prototype-based guidance.
In this paper, we leverage whole slide images at two resolution levels - low magnification (5x) and high magnification (10x) - to capture both tissue-level context and fine cellular detail. TThe whole slide images are processed by Otsu’s algorithm to remove all the non-tissue regions. The cropped patch size is $256\times 256$. All experiments were executed on a single NVIDIA A100 GPU node.